OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 6)(x − 6) and (x − 6)(x − 6) Help Please?

OpenStudy (anonymous):

x^2 - 12 and x^2 - 12x + 12 but i dont get how you mean to write a sentence .....

OpenStudy (anonymous):

you're wrong @Sierra55555

OpenStudy (anonymous):

no im not .

OpenStudy (anonymous):

then you answer it

OpenStudy (anonymous):

oh i added the 6's im sorry ! it would be 36's at the ends instead of that !

OpenStudy (anonymous):

(x + 6)(x − 6) = x(x-6) + 6(x-6) (x+6)(x+6) = x(x+6) +x(x+6) distribute it and see if you can see the difference

OpenStudy (anonymous):

crap i screwed up the 2nd equation (x+6)(x+6) = x(x+6) +6(x+6)